# The effect of control barrier functions on energy transfers in controlled physical systems

### v1

@@ III-B Beyond damping injection
Consider the case $R=0$, that is, the system (1) evolves on energetic isolines.
Using a CBF $h(x)=c-H(x)$, i.e., $\mathcal{C}=\{x\in\mathcal{D}|H(x)\leq c\}$, the power equality (8) becomes:
$\dot{H}=P_{h,\alpha}=-\mathbbm{1}_{\Psi<0}(\alpha(c-H(x)),\qquad\Psi=\alpha(c-H(x)).$ (10)
@@
If $x\in\mathcal{C}$ then $P_{h,\alpha}=0$, while if $x\notin\mathcal{C}$ the safety-critical control extracts the negative power $P_{h}=\alpha(h)$, which can be modulated with $\alpha$. In case the $\alpha$ is linear, i.e., $\alpha(h)=\gamma h,\,\gamma\in\mathbb{R}^{+}$, the constant $\gamma$ acts as a proportional gain on the “energy error” $c-H(x)$ measuring the distance from the closest energy value that the system is ultmately allowed to stay in. The (simple) stability of the minimum of the Hamiltonian is clearly preserved as $[h,H]_{gg^{\top}}=-[H,H]_{gg^{\top}}\leq 0$.
Using instead $h(x)=-c+H(x)$, i.e., $\mathcal{C}=\{x\in\mathcal{D}|H(x)\geq c\}$, the power equality (8) becomes:
$\dot{H}=P_{h,\alpha}=\mathbbm{1}_{\Psi<0}(\alpha(c-H(x)),\qquad\Psi=\alpha(H(x)-c).$ (11)
@@
If $x\in\mathcal{C}$ then $P_{h,\alpha}=0$, but now if $x\notin\mathcal{C}$, $P_{h,\alpha}$ is positive, corresponding to a non-passive action, as the safety-critical controller needs to inject power into the system to reach the set $\mathcal{C}$, which can be again modulated by $\alpha$.
@@ IV Mechanical systems and energy-based CBFs
In order to better comprehend the proposed methodology and discuss how it complements standard energy-based approaches, let us specialise system (1) to fully actuated mechanical systems, where we introduce the state $x=(q^{\top},p^{\top})^{\top}\in\mathbb{R}^{2n}$ as canonical Hamiltonian coordinates on the cotangent bundle of the configuration manifold of the system. In particular $q\in\mathbb{R}^{n}$ is the vector of generalized coordinates. $p\in\mathbb{R}^{n}$ denotes the generalized momenta, $p:=M(q)\dot{q}$, where $M(q)=M(q)^{\top}>0$ is the positive definite inertia matrix of the system. The equations of motions in canonical form are given by (1) with
@@
where $H:\mathbb{R}^{2n}\to\mathbb{R}$ is the total energy (Hamiltonian)
@@
where $K_{e}(q,p)=\frac{1}{2}p^{\top}M^{-1}(q)p$ is the kinetic energy, $V:\mathbb{R}^{n}\to\mathbb{R}$ maps the position state to conservative potentials (gravity, elastic effects), $D=D^{\top}\geq 0$ takes into account dissipative and friction effects, $B\in\mathbb{R}^{n\times n}$ is the (full rank) input matrix [^2] For lighten notation we hide possible state dependencies on $D$ and $B$., $I_{n}$ indicates the $n\times n$ identity matrix and non specified dimensions of matrices, comprising those with only zero entries and denoted with the symbol “00”, are clear from the context.
@@
$P_{h,\alpha}=-\mathbbm{1}_{\Psi<0}\frac{\dot{q}^{\top}BB^{\top}\frac{\partial h}{\partial p}}{\frac{\partial^{\top}h}{\partial p}BB^{\top}\frac{\partial h}{\partial p}}\Psi,$ (14)
@@
The previous expressions allow to design safety-critical controllers with a qualitiative and quantitative insight on the effect of a certain CBF $h(q,p)$ on the power balance of the controlled mechanical system. An interesting class of CBFs which deserves some discussion is
@@
where $\bar{h}$ is a smooth function of the configuration variable only and $c$ is a real number.
The sufficient condition of Proposition 1 on stability-preserving safety-critical control (i.e., the condition on which safety-critical control always injects damping in the controlled mechanical system) reduces to
$\mathbbm{1}_{\Psi<0}\bigg{(}\dot{q}^{\top}BB^{\top}\frac{\partial h}{\partial p}\bigg{)}\leq 0$ (17)
@@
CBFs (18) encompass those referred to as energy-based CBFs in [Singletary2021Safety-CriticalSystems], defined as (18) with $c=0$ and $\bar{h}=\alpha_{E}h_{k}(q)$ where $h_{k}(q)$ is a function whose superlevel safe set represents purely kinematic constraints (e.g., obstacle avoidance). In [10136379] the safety-critical control induced by CBFs in the form (18) was proven to preserve passivity of nominal controllers, similarly to the discussed in Proposition 1, where $H$ is assumed to be a Lyapunov function.
@@
This result provides further interpretability in the structure of $\Psi$, whose negativity determines when the safety-critical control starts to act, and whose absolute value determines the amount of power which is injected (if (16) is taken with the plus) or dissipated (if (16) is taken with the minus) instantaneously in the system. We give practical insight with the following examples.
Example 2 (Example 1 applied to mechanical systems).
Notice that for CBFs of the form (18) (pure damping injection), the first case in Example 1 is retrieved if we use $\bar{h}=-V$, and thus $\{h,H\}=0$ and the value of $\Psi$ reduces to (10) if friction is ignored. Similarly, one can retrieve the second case in Example 1, i.e., choosing a CBF (16) with the plus sign and $\bar{h}=V$.
The latter example is a particular case in which the positivity of $\Psi$ depends only on whether the system is in the safe set, but more in general the result in Proposition 2 allows one to get information about the way safety-critical control acts in terms of energy transfer and use that information in the CBF design. As an example, we present a case of safety-critical control that involves the prevention of several physical safety hazards in robotics by limiting the total kinetic energy of the system.
Example 3 (Limiting kinetic energy).
If one wants to limit kinetic energy in a mechanical system (12), the candidate CBF is $h(q,p)=-K_{e}(q,p)+c$, a particular case of (18) with $\bar{h}=0$. Notice that in this case, if there is no friction, we obtain $\Psi=\dot{V}+\alpha(h(q,p))$, which encodes the information on when the safety-critical effects occur, how much damping is injected in the system, and how the design of the class $\mathcal{K}$ function $\alpha$ could be used to influence the resulting dynamics.
V Simulations
Figure 1: Mass-Spring system. Bounding energy from above (left) and below (right). From above: i) phase space trajectory, ii) CBF, iii) safety-critical control input, and iv) total energy.
Figure 2: Mass-Spring system. Limiting kinetic energy starting from inside (left) and outside (right) the safe set. From above: i) phase space trajectory, ii) CBF, iii) safety-critical control input, and iv) total energy.
Figure 3: Double pendulum. Limiting kinetic energy starting from inside (left) and outside (right) the safe set. From above: i) CBF, ii) safety-critical control input, and iii) total energy.
In this section, we display and discuss numerical simulations of mechanical systems undergoing safety-criticl control with CBFs in the form (16), previously discussed in Examples 2 and 3. When the phase space is plotted, the safe set is represented in white and its complement is represented in gray. The class $\mathcal{K}$ function $\alpha$ is chosen linearly, i.e., $\alpha(h)=\gamma h$, and simulations are performed for different positive scalars $\gamma$.
@@ V-1 Bounding total energy from below and above
To display the validity and the potential of the proposed results we initially consider as a plant a simple oscillator (mass-spring) with no dissipation, i.e., system (12) with $D=0$, $B=1$, $H=\frac{p^{2}}{2m}+\frac{kq^{2}}{2}$, $k=0.5$ and $m=2$. The results are shown in Fig. 1. The left column implements a safety-critical controller given by a CBF bounding the total energy from above ($h=-H+c$) while the second column uses a CBF bounding the total energy from below ($h=H-c$), with energy limit $c=10$. It is interesting to see how the proposed approach generates new ways to implement mechanical oscillations at the desired energy level. As theoretically derived from example 1, we see that the safety critical input is triggered once the system is outside the safe set, and we see how the $\gamma$ parameter can be set to regulate the amount of power the controller dissipates or injects.
@@ V-2 Bounding kinetic energy
Fig. 2 shows simulations with the same plant as before, but with a CBF designed to limit only the kinetic energy, i.e., $h=-K_{e}+c$ where $c=20$, as explained in Example 3. The left column corresponds to experiments in which the system starts in the safe set, and the right column to one in which the system starts outside the safe set. These simulations clearly show both the invariance and the asymptotic stability of the safe set. Looking at the total energy, it appears clearly that, as predicted by the theory, the safety-critical controller acts purely as a damper, with the amount of damping which be regulated with $\gamma$. As explained in Example 3, the activation of safety-critical effects is no longer trivially related to whether the system is in the safe set or not, and can be inspected by checking the positivity of the constraint (20).
Finally, in Fig. 3 we show simulations for a double pendulum with masses $m_{1}=m_{2}=2$ and lengths $l_{1}=l_{2}=1$, without natural dissipation, in a gravitational field $V(q)$ (the zero potential corresponds to the reference configuration in which the pendulum is in the horizontal position). This system is described by (12) with $D=0$, $B=I_{2}$ and $H=\frac{p^{\top}M^{-1}(q)p}{2}+V(q)$. The CBF is designed to limit only the kinetic energy, i.e., $h=-K_{e}+c$ where $c=10$. For this system, the natural dynamics are in general chaotic for sufficiently high levels of initial energy, and the presence of safety-critical effects can be analyzed by inspecting, as before, the positivity of (20). As clear from the plots and predicted by the theory, the controller act purely as a damper (the total energy is non increasing), preserving simple stability of the origin of the phase space. The system converges to solutions with the desired level of energy, which depends both on the choice of $c$ and $\gamma$.
@@ VI Conclusions
In this paper we gave constructive tools to study the qualitative and quantitative effect of safety-critical control schemes implemented with CBFs on the energy balance of controlled physical systems. The analysis led to novel energy-aware schemes, such as selective damping injection mechanisms and active control strategies that inject energy into the controlled system to achieve desired closed-loop behaviors. We are working towards experimental implementations of the proposed algorithms.

### v2

@@ III-2 CBF-based energy-aware designs
Consider the case $R=0$, that is, the system (1) evolves on energetic isolines. Using a CBF $h(x)=c-H(x)$, i.e., $\mathcal{C}=\{x\in\mathcal{D}|H(x)\leq c\}$, the power equality (8) becomes:
$\dot{H}=P_{h,\alpha}=-\mathbb{I}_{\Psi<0}(\alpha(c-H(x)),\qquad\Psi=\alpha(c-H(x)).$ (10)
@@
If $x\in\mathcal{C}$ then $P_{h,\alpha}=0$, while if $x\notin\mathcal{C}$ the safety-critical control extracts the negative power $P_{h}=\alpha(h)$, which can be modulated with $\alpha$. In case the $\alpha$ is linear, i.e., $\alpha(h)=\gamma h,\,\gamma\in\mathbb{R}^{+}$, the constant $\gamma$ acts as a proportional gain on the “energy error" $c-H(x)$ measuring the distance from the closest energy value that the system is ultmately allowed to stay in. The (simple) stability of the minimum of the Hamiltonian is clearly preserved as $[h,H]_{gg^{\top}}=-[H,H]_{gg^{\top}}\leq 0$. Using instead $h(x)=-c+H(x)$, i.e., $\mathcal{C}=\{x\in\mathcal{D}|H(x)\geq c\}$, the power equality (8) becomes:
$\dot{H}=P_{h,\alpha}=\mathbb{I}_{\Psi<0}(\alpha(c-H(x)),\qquad\Psi=\alpha(H(x)-c).$ (11)
@@
If $x\in\mathcal{C}$ then $P_{h,\alpha}=0$. If instead $x\notin\mathcal{C}$, $P_{h,\alpha}$ is positive, i.e., the safety-critical controller injects power (modulated by $\alpha$) into the system to reach $\mathcal{C}$. The system is shaped into a nonlinear oscillator at the desired energy level, a behavior achieved by a non passive control action.
@@ IV Mechanical systems and energy-based CBFs
In order to better comprehend the proposed methodology and discuss how it complements standard energy-based approaches, let us specialise system (1) to fully actuated mechanical systems, where we introduce the state $x=(q^{\top},p^{\top})^{\top}\in\mathbb{R}^{2n}$ as canonical Hamiltonian coordinates on the cotangent bundle of the configuration manifold of the system. In particular $q\in\mathbb{R}^{n}$ is the vector of generalized coordinates, $p\in\mathbb{R}^{n}$ denotes the generalized momenta, $p:=M(q)\dot{q}$, where $M(q)=M(q)^{\top}>0$ is the positive definite inertia matrix of the system. The equations of motions are given by (1) with
@@
where $H:\mathbb{R}^{2n}\to\mathbb{R}$ is the total mechanical energy
@@
where $K_{e}(q,p)=\frac{1}{2}p^{\top}M^{-1}(q)p$ is the kinetic energy, $V:\mathbb{R}^{n}\to\mathbb{R}$ maps the position state to conservative potentials (gravity, elastic effects), $D=D^{\top}\geq 0$ takes into account dissipative and friction effects, $B\in\mathbb{R}^{n\times n}$ is the (full rank) input matrix, $I_{n}$ indicates the $n\times n$ identity matrix and non specified dimensions of matrices are clear from the context.
@@
$P_{h,\alpha}=-\mathbb{I}_{\Psi<0}\frac{\dot{q}^{\top}BB^{\top}\frac{\partial h}{\partial p}}{\frac{\partial^{\top}h}{\partial p}BB^{\top}\frac{\partial h}{\partial p}}\Psi,$ (14)
@@
The previous expressions allow to design safety-critical controllers with a qualitiative and quantitative insight on the effect of a certain CBF $h(q,p)$ on the power balance of the controlled mechanical system. A class of CBFs which deserves some discussion is
@@
where $\bar{h}$ is any differentiable function of the configuration variable and $c$ is a real number.
The sufficient condition of Proposition 1 on stability-preserving safety-critical control (i.e., the condition on which safety-critical control acts as a damper) reduces to
$\mathbb{I}_{\Psi<0}\bigg{(}\dot{q}^{\top}BB^{\top}\frac{\partial h}{\partial p}\bigg{)}\leq 0.$ (17)
@@
CBFs (18) encompass those referred as energy-based CBFs in [Singletary2021Safety-CriticalSystems], defined as (18) with $c=0$ and $\bar{h}=\alpha_{E}h_{k}(q)$ where $h_{k}(q)$ is a CBF encoding purely kinematic constraints (e.g., obstacle avoidance). In [10136379] the safety-critical control induced by CBFs in the form (18) was proven to preserve passivity of nominal controllers.
@@
Expression (20) provides interpretability in the structure of $\Psi$ and its dependence on the CBF parameter $\alpha$. A negative $\Psi$ indicates that the safety-critical controller is active, and (20) is the injected or extracted power (depending on the sign of (16)).
V Simulations and Discussion
Figure 1: Mass-Spring system. Bounding energy from above (left) and below (right). From above: i) phase space trajectory, ii) CBF, iii) control input, and iv) total energy.
Figure 2: Limiting kinetic energy from inside (left) and outside (right) the safe set. From above: i) phase space trajectory, ii) CBF, iii) control input, and iv) total energy.
In this section, we display and discuss numerical simulations of mechanical systems undergoing safety-criticl control with CBFs in the form (16). When the phase space is plotted, the safe set is represented in white and its complement is represented in gray. The function $\alpha$ is chosen as $\alpha(h)=\gamma h$, and simulations are performed for different positive scalars $\gamma$. When not reported, units are clear from the context.
@@ V-1 Bounding total energy from below and above
We initially consider as a plant a simple oscillator (mass-spring) with no dissipation, i.e., system (12) with $D=0$, $B=1$, $H=\frac{p^{2}}{2m}+\frac{kq^{2}}{2}$, $k=0.5$ and $m=2$. The results are shown in Fig. 1. The left column implements a safety-critical controller given by a CBF bounding the total energy from above ($h=-H+c$) while the second column uses a CBF bounding the total energy from below ($h=H-c$), with energy limit $c=10$. The proposed approach generates new ways to implement mechanical oscillations at the desired energy level. As theoretically derived from example 1, we see that the safety critical input is triggered once the system is outside the safe set, and we see how the $\gamma$ parameter can be set to regulate the power the controller dissipates or injects.
@@ V-2 Bounding kinetic energy
Fig. 2 shows simulations with the same plant, but with a CBF designed to limit only the kinetic energy, i.e., $h=-K_{e}+c$ where $c=20$. The left column corresponds to experiments in which the system starts in the safe set, and the right column to one in which the system starts outside the safe set. The results show both the invariance and the asymptotic stability of the safe set. Looking at the total energy, it appears clearly that, as predicted by the theory, the safety-critical controller acts purely as a damper, with the amount of damping which be regulated with $\gamma$. The activation of safety-critical effects is no longer trivially related to whether the system is in the safe set or not, and can be inspected by checking the positivity of the constraint (20).
V-3 Energy-based nonlinear oscillator
In Fig. 3 we show simulations for a nonlinear plant: a double pendulum with masses $m_{1}=m_{2}=1.5$ and lengths $l_{1}=l_{2}=1$, with natural dissipation $b=0.3$ in both the joints. This system is described by (12) with $D=bI_{2}$, $B=I_{2}$ and $H=\frac{p^{\top}M^{-1}(q)p}{2}+V(q)$, with $V(q)$ the gravitational potential, taking negative values below the first joint. The CBF is designed to limit the total energy, i.e., $h=-H+c$ where $c=-40$, and it is activated after $10s$ from the start of the simulation. The system, which in the first phase was dissipating its initial mechanical energy, converges to the safe set after the CBF activation. Interestingly, the controlled system converges to a limit cycle: the natural dissipation and the power injected by the CBF counterbalance each other on the time period on which the limit cycle emerges, resembling a periodic locomotion task. This procedure constitutes a new, energy-based, way to generate periodic motions in mechanical systems.
Figure 3: Double pendulum. CBF imposing total energy lower limit is activated after 10 $\mathrm{s}$. From above, left column: i) CBF, ii) safety-critical control input, iii) total energy, and iv) power terms in (13), detailed on its right. Top right: cartesian snapshots of the system before (gray) and after (black) the CBF activation, and joint trajectory below.
@@ VI Conclusions
In this paper we gave constructive tools to study the qualitative and quantitative effect of safety-critical control schemes implemented with CBFs on the energy balance of controlled physical systems. The analysis led to novel energy-aware schemes, such as selective damping injection mechanisms and active control strategies that inject energy into the controlled system to achieve desired closed-loop behaviors.